For MAC= 36-2E and MD= E,
(i) What is the socially desirable level of emissions?
(ii) What is the amount of net social benefits received when the emission level is reduced to 10 units (i.e. E=10)?
Answer #1

Part 1 ) At socially desirable level of emissions, MAC = MD

Using MAC= 36-2E and MD= E,

MAC = MD

= 36-2E = E

= 36 = 3E

E = 12

Socially desired level = 12

Part 2) When E = 10,

MD = E = 10

MAC = 36-2E = 36-2*10 = 16

Net Social Benefits = MAC - MD = 16 – 10 = 6